In the lesson linear inequalities ( solution of system) how to decide the values of X and Y axes on the graph in order to draw the slope

Grade:11

1 Answers

Arun
5 years ago
Put the line in the form 
 
x/a + y/b  = 1
Remember RHS should be 1.
For example
2x + y = 4
Hence divide the equation by 4
2x/4 + y/4 = 1
x/2 + y/4 = 1
Hence on x axis (2, 0) and on y axis (0,4 ) are plot